# Nonlinear Concept Erasure: a Density Matching Approach

### v1

@@ Abstract
Ensuring that neural models used in real-world applications cannot infer sensitive information, such as demographic attributes like gender or race, from text representations is a critical challenge when fairness is a concern. We address this issue through concept erasure, a process that removes information related to a specific concept from distributed representations while preserving as much of the remaining semantic information as possible. Our approach involves learning an orthogonal projection in the embedding space, designed to make the class-conditional feature distributions of the discrete concept to erase indistinguishable after projection. By adjusting the rank of the projector, we control the extent of information removal, while its orthogonality ensures strict preservation of the local structure of the embeddings. Our method, termed $\overline{\mathrm{L}}$, achieves state-of-the-art performance in nonlinear erasure of a discrete attribute on classic natural language processing benchmarks. Furthermore, we demonstrate that $\overline{\mathrm{L}}$ effectively mitigates bias in deep nonlinear classifiers, thereby promoting fairness.
@@ 1 Introduction
Figure 1: $\overline{\mathrm{L}}$ erases a discrete attribute $z$ by learning an orthogonal projection $P$ onto a subspace $\mathcal{V}^{\perp}$ of the feature space in which the distributions of the projected samples $x^{\perp}=Px$ are indistinguishable across values of the protected attribute to erase $z$.
@@
We propose a simple yet novel concept erasure method which learns an orthogonal projection in the embedding space. The optimization process is driven by a density-matching erasure objective, ensuring that the class-conditional distributions of the representations after erasure are indistinguishable. The projection rank can be adjusted to control the extent of information removal. Leveraging orthogonal projections satisfies the dual objective of achieving non-invertible erasure while preserving the local structure of the feature space. We refer to our framework as $\overline{\mathrm{\textbf{L}}}$: Non-Linear concept Erasure via Orthogonal Projection and Alignment of post-erasure Representation Distributions. To evaluate the effectiveness of our approach, we conducted experiments demonstrating state-of-the-art performance in discrete concept erasure for the task of demographic attribute removal on standard NLP benchmarks. Furthermore, we show empirically that $\overline{\mathrm{L}}$ can improve fairness of downstream classifiers with moderate impact on their performance. Our contributions are:
  • We introduce $\overline{\mathrm{L}}$, a novel method for unconstrained nonlinear concept erasure relying on orthogonal projections and a class-conditional density matching objective.
  • We demonstrate that $\overline{\mathrm{L}}$ achieves state-of-the-art performance on many NLP benchmarks for the task of demographic attributes erasure.
  • We demonstrate the utility of $\overline{\mathrm{L}}$ in improving the fairness of deep classifiers with a moderate impact on performance.
@@ 2 Related work
Constrained de-biasing, or erasure, refers to optimizing post-erasure representations for a predefined downstream task, such as a classification [3], or a more generic one, such as Masked Language Modeling (MLM) for language models fine-tuning [11, 13]. In contrast, downstream tasks are assumed unknown a priori in an unconstrained mode, making it essential to preserve semantic information for broad applicability. Semantic preservation is typically evaluated via semantic similarity tests (e.g., WordSim-353 [2], WHEAT [9]) or the preservation of the neighborood structure [4]. This paper is a contribution to unconstrained erasure.
@@ 3 Learning Projections with a Density Matching Objective
This section presents our optimization method, $\overline{\mathrm{L}}$, which aims to achieve nonlinear erasure of a discrete concept from a set of representations through orthogonal projections. We first introduce a parametrization for learning an orthogonal projection of prescribed rank, and subsequently introduce a density matching criterion to select the projection that best removes the concept information.
@@ 3.1 Learning Orthogonal Projections
Our approach follows the linear bias subspace hypothesis, which posits that, for a given set of representations, the information related to a specific concept is entirely contained within a subspace of the embedding space [7, 27]. Consequently, we seek an orthogonal decomposition of the embedding space $\mathbb{R}^{d}$ as $\mathbb{R}^{d}=\mathcal{V}^{\perp}\oplus\mathcal{V}^{\parallel}$ where $\mathcal{V}^{\perp}$ is the subspace of dimension $r$ that is to be preserved, and $\mathcal{V}^{\parallel}$ is its orthogonal complement of dimension $d-r$ that is to be neutralized.
@@ 3.2 Class-Conditional Density Matching
[…]
@@
Algorithm 1 $\overline{\mathrm{L}}$
@@
7:Project onto the closest orthogonal projector to $UU^{\top}$ in Frobenius norm
8:$U,\Lambda=\text{EigenDecomposition}(UU^{\top})$
@@
10:$P\leftarrow U_{r}{U_{r}}^{\top}$
Algorithm 2 Cascaded $\overline{\mathrm{L}}$
@@
5:Erase the concept via $\overline{\mathrm{L}}$ using Algorithm 1:
6:$P^{\prime}=\overline{\mathrm{L}}\text{EOPARD}(\{U_{L}^{\top}x_{i}\}_{i=1}^{N},\mathcal{C},r,\gamma,U^{(0)},N_{e})$
@@
3.3 $\overline{\mathrm{\textbf{L}}}$
The optimization procedure for $\overline{\mathrm{L}}$ is outlined in Algorithm 1. The goal is to learn a rank-$r$ orthogonal projection $P$ that erases a discrete concept nonlinearly, parameterized via a matrix $U\in\mathbb{R}^{d\times r}$ such that $P=UU^{\top}$, as in Eq. 2. The learning objective to minimize during training combines the projection loss $\mathcal{L}_{p}$ (Eq. 3) and the erasure loss $\mathcal{L}_{\mathrm{MMD}}$ (Eq. 6) and is given by:
@@ 3.3 $\overline{\mathrm{\textbf{L}}}$
The number of trainable parameters in $\overline{\mathrm{L}}$ is $d\cdot r$, significantly lower than that of existing nonlinear erasure methods such as FaRM [3] and KRaM [4], which involve training multi-layer perceptrons for concept removal.
In $\overline{\mathrm{L}}$, the adjustable rank $r$ governs the trade-off between the preservation of information and the effectiveness of concept removal. There does not exist a universally optimal value for $r$; the rank adjustability constitutes a key strength of our method, enabling trade-offs between erasure and information preservation that are not possible with more rigid approaches.
Cascaded $\overline{\mathrm{\textbf{L}}}$
@@
Cascaded erasure is a two-stage process that integrates linear and nonlinear erasure. In the first stage, we apply an orthogonalized LEACE projection [5] to linearly remove the concept. The embeddings are then projected onto the $(d-K+1)$-dimensional image subspace of this projection. In the second stage, these linearly "sanitized" embeddings are further processed using a nonlinear erasure method. When instantiated with $\overline{\mathrm{L}}$, it amounts to computing two orthogonal projections acting on nested subspaces. The overall cascaded projection remains orthogonal and inherits the benefits of both linear and nonlinear concept erasure. The full procedure for Cascaded $\overline{\mathrm{L}}$ is outlined in Algorithm 2.
@@ 4.1 Datasets
is a real world dataset suited for studying gender biases in biography classification tasks [10]. It consists in biographies collected through web scraping and labeled with binary gender and occupation (28 in total). We use the version from [23], which contains $\sim$98of the original dataset, as the full version is no longer publicly available. It consists of 399,423 biographies stratified by occupation, split 65%/10%/25% across training, validation, and test.
@@ 4.2 Practical Aspects of MMD Evaluation
An important issue when using the MMD estimator with a Gaussian kernel in practice is the selection of the kernel bandwidth. To mitigate the necessity of tuning $\sigma$ for each experiment, it is usual to rely on a heuristic [16]. In this study, we rely on the mean heuristic which consists in setting $\sigma$ to be the pairwise mean distance on the aggregate sample, i.e. $\sigma^{2}=\frac{\sum_{x,x^{\prime}}||x-x^{\prime}||^{2}}{N^{2}-N}$.
@@ 4.3 Evaluation metrics
When there is no predefinite downstream task for which we can evaluate the accuracy, to measure the overall preservation of the semantic information in the embedding space we rely on the metric $A_{\mathbf{k}}$ introduced in [4] that quantifies the average overlap between the $k$-nearest neighbor sets of $x_{i}$ and $x_{i}^{\perp}$. Formally,
$A_{k}=\frac{1}{N}\sum_{x\in\mathcal{D}}\left[\frac{1}{k}|\mathrm{knn}(x)\cap\mathrm{knn}(x^{\perp})|\right]\,,$ (9)
where $\mathrm{knn}(\cdot)$ computes the $k$-nearest neighbor set of a representation. The metric ranges from 00 to $1$, with values closer to $1$ indicating greater preservation of the local geometry of the embeddings. We set $k=\frac{|\mathcal{D}|}{2}$, denoting the average neighborhood overlap as $A_{50\%}$.
@@
$\mathrm{DP}=\sum_{y\in\mathcal{Y}}|p(\hat{Y}=y|Z=c)-p(\hat{Y}=y|Z=c^{\prime})|\$ (11)
@@ 4.4 Implementation Details
All models are trained using the Adam optimizer [17]. For $\overline{\mathrm{L}}$, weight decay is set to zero to avoid interfering with the optimization of the projection loss (Eq. 3).
@@ 5.1 Nonlinear Concept Erasure
$\overline{\mathrm{L}}$ enables effective erasure of information associated with a discrete concept, with the projector rank $r$ serving as a tunable parameter that controls the strength of the erasure. Empirically, we observe that the difficulty of recovering the concept post-erasure increases steadily as the number of preserved dimensions decreases (see Figure 3, left panels). In Table 1, we exhibit configurations where $\overline{\mathrm{L}}$ outperforms nonlinear erasure baselines in terms of MDL or probing accuracy which drops to near-random levels. Notably, in Table 1(a), gender treated as a ternary concept during training and evaluation is almost fully erased, demonstrating the validity of $\overline{\mathrm{L}}$ beyond binary settings. Qualitatively, the vizualizations in figure 2 and in figures 6,7 in Supplementary Material D illustrate that concept class groups, which were previously distinguishable, become indistinguishable following erasure via $\overline{\mathrm{L}}$.
Figure 2: t-SNE visualization of GloVe embeddings before (left) and after (right) erasure via Cascaded $\overline{\mathrm{L}}$ with a rank-150 orthogonal projection. Gender is a ternary concept. Male-biased, female-biased, and neutral words are displayed in different colors.
Results reported in the left panels of Figure 3 highlights the superior erasure capabilities of Cascaded $\overline{\mathrm{L}}$ (Algorithm 2). For a fixed projection rank, the difficulty of recovering the concept post-erasure is consistently higher with Cascaded $\overline{\mathrm{L}}$ compared to the standard training (Algorithm 1). For a given value of MDL, Cascaded $\overline{\mathrm{L}}$ retains a larger number of dimensions in the feature space. This confirms our intuition that $\overline{\mathrm{L}}$ trained in a standard way struggles with the removal of strong linear signals, which undermines nonlinear erasure performance. However, concept erasure is achieved with standard $\overline{\mathrm{L}}$ at low rank. Both FaRM and KRaM, in their standard forms, perform well in terms of both linear and nonlinear erasure, suggesting that implementing them in a cascaded setup should not improve their performance, which we confirm experimentally (see Supplementary Material D).
[…]
@@
$a_{z}^{\mathrm{bin}}$downarrow MDL $\uparrow$ $a_{z}^{\mathrm{ter}}$downarrow $A_{50\%}\uparrow$ WS-353 $\uparrow$
@@
$a_{z}$downarrow MDL $\uparrow$ $a_{y}$uparrow
@@
$a_{z}$downarrow MDL $\uparrow$ $a_{y}$uparrow
@@
The right panels of Figure 3 highlight a consistent trade-off between concept erasure and the utility of post-erasure representations across all benchmark datasets, with optimal configurations located in the upper-right regions. The rank $r$, which controls the strength of erasure, modulates this trade-off. FaRM and KRaM achieve strong erasure but at the cost of substantial degradation in utility for gender, as evidenced by their positions in the upper-left regions of the left panels in Figures 3a and 3b.
The adjustable rank of $\overline{\mathrm{L}}$ offers flexibility in navigating this trade-off, enabling configurations that outperform baselines across several settings. Notably, on GloVe embeddings, $\overline{\mathrm{L}}$ better preserves both local neighborhood structure and semantic content (Table 1(a)).
On Bias in Bios, FaRM and KRaM outperform $\overline{\mathrm{L}}$ in terms of MDL-based erasure at the expense of downstream task performance. At low rank, $\overline{\mathrm{L}}$ achieves trade-offs between probing accuracy and downstream task accuracy comparable to baselines (e.g., $r=10$ and $r=50$ for standard and cascaded training, respectively, in Table 1(b)). Figure 3(b) reveals that beyond a certain threshold, performance plateaus; further reducing $r$ yields marginal gains in erasure while significantly harming utility. In such cases, $\overline{\mathrm{L}}$ (e.g., at $r=150$ with cascaded training) allows for a more satisfying balance between the reduction of the probing accuracy and downstream accuracy compared to baselines.
On DIAL, for race erasure, $\overline{\mathrm{L}}$ remains competitive without outperforming the baselines. Relying on linear projections in $\overline{\mathrm{L}}$ provides advantages in simplicity, interpretability, and parameter efficiency. However, such linear methods are inherently limited when addressing the erasure of concepts encoded nonlinearly. While erasing even a small number of dimensions can significantly reduce the recoverability of the target concept, achieving strong erasure often requires neutralizing a large subspace, which may degrade the utility of the resulting representations.

### v2

@@ Abstract
Ensuring that neural models used in real-world applications cannot infer sensitive information, such as demographic attributes like gender or race, from text representations is a critical challenge when fairness is a concern. We address this issue through concept erasure, a process that removes information related to a specific concept from distributed representations while preserving as much of the remaining semantic information as possible. Our approach involves learning an orthogonal projection in the embedding space, designed to make the class-conditional feature distributions of the discrete concept to erase indistinguishable after projection. By adjusting the rank of the projector, we control the extent of information removal, while its orthogonality ensures strict preservation of the local structure of the embeddings. Our method, termed $\overline{\mathrm{L}}$EOPARD, achieves state-of-the-art performance in nonlinear erasure of a discrete attribute on classic natural language processing benchmarks. Furthermore, we demonstrate that $\overline{\mathrm{L}}$EOPARD effectively mitigates bias in deep nonlinear classifiers, thereby promoting fairness.
@@ 1 Introduction
Figure 1: $\overline{\mathrm{L}}$EOPARD erases a discrete attribute $z$ by learning an orthogonal projection $P$ onto a subspace $\mathcal{V}^{\perp}$ of the feature space in which the distributions of the projected samples $x^{\perp}=Px$ are indistinguishable across values of the protected attribute to erase $z$.
@@
We propose a simple yet novel concept erasure method which learns an orthogonal projection in the embedding space. The optimization process is driven by a density-matching erasure objective, ensuring that the class-conditional distributions of the representations after erasure are indistinguishable. The projection rank can be adjusted to control the extent of information removal. Leveraging orthogonal projections satisfies the dual objective of achieving non-invertible erasure while preserving the local structure of the feature space. We refer to our framework as $\overline{\mathrm{\textbf{L}}}$EOPARD: Non-Linear concept Erasure via Orthogonal Projection and Alignment of post-erasure Representation Distributions. To evaluate the effectiveness of our approach, we conducted experiments demonstrating state-of-the-art performance in discrete concept erasure for the task of demographic attribute removal on standard NLP benchmarks. Furthermore, we show empirically that $\overline{\mathrm{L}}$EOPARD can improve fairness of downstream classifiers with moderate impact on their performance. Our contributions are:
  • We introduce $\overline{\mathrm{L}}$EOPARD, a novel method for unconstrained nonlinear concept erasure relying on orthogonal projections and a class-conditional density matching objective.
  • We demonstrate that $\overline{\mathrm{L}}$EOPARD achieves state-of-the-art performance on many NLP benchmarks for the task of demographic attributes erasure.
  • We demonstrate the utility of $\overline{\mathrm{L}}$EOPARD in improving the fairness of deep classifiers with a moderate impact on performance.
@@ 2 Related work
Constrained de-biasing, or erasure, refers to optimizing post-erasure representations for a predefined downstream task, such as a classification [3], or a more generic one, such as Masked Language Modeling (MLM) for language models fine-tuning [11, 13]. In contrast, downstream tasks are assumed unknown a priori in an unconstrained mode, making it essential to preserve semantic information for broad applicability. Semantic preservation is typically evaluated via semantic similarity tests (e.g., WordSim-353 [2], WHEAT [9]) or the preservation of the neighborhood structure [4]. This paper is a contribution to unconstrained erasure.
@@ 3 Learning Projections with a Density Matching Objective
This section presents our optimization method, $\overline{\mathrm{L}}$EOPARD, which aims to achieve nonlinear erasure of a discrete concept from a set of representations through orthogonal projections. We first introduce a parametrization for learning an orthogonal projection of prescribed rank, and subsequently introduce a density matching criterion to select the projection that best removes the concept information.
@@ 3.1 Learning Orthogonal Projections
Our approach follows the linear bias subspace hypothesis, which posits that, for a given set of representations, the information related to a specific concept is entirely contained within a subspace of the embedding space [7, 27]. Consequently, we seek an orthogonal decomposition of the embedding space $\mathbb{R}^{d}$ as $\mathbb{R}^{d}=\mathcal{V}^{\perp}\oplus\mathcal{V}^{\parallel}$ where $\mathcal{V}^{\perp}$ is the subspace of dimension $r$ that is to be preserved, and $\mathcal{V}^{\parallel}$ is its orthogonal complement of dimension drd-r that is to be neutralized.
@@ 3.2 Class-Conditional Density Matching
$\begin{aligned} \widehat{\mathrm{MMD}^{2}}(\mathbb{P}_{z=1}^{\perp},\mathbb{P}_{z=2}^{\perp})&=\frac{1}{n_{z=1}(n_{z=1}-1)}\sum_{\begin{subarray}{c}x,x^{\prime}\in\mathcal{D}_{z=1}\\ x\neq x^{\prime}\end{subarray}}k(x^{\perp},x^{\prime\perp})\\ &\quad+\frac{1}{n_{z=2}(n_{z=2}-1)}\sum_{\begin{subarray}{c}x,x^{\prime}\in\mathcal{D}_{z=2}\\ x\neq x^{\prime}\end{subarray}}k(x^{\perp},x^{\prime\perp})\\ &\quad-\frac{2}{n_{z=1}n_{z=2}}\sum_{x\in\mathcal{D}_{z=1}}\sum_{x^{\prime}\in\mathcal{D}_{z=2}}k(x^{\perp},x^{\prime\perp}).\end{aligned}$
@@
Algorithm 1 $\overline{\mathrm{L}}$EOPARD
@@
7:% Project onto the closest orthogonal projector to $UU^{\top}$ in Frobenius norm
@@
9:UrU[$_{r}\leftarrow U[:,1{:}$r]
[…]
Algorithm 2 Cascaded $\overline{\mathrm{L}}$EOPARD
@@
5:Erase the concept via $\overline{\mathrm{L}}$EOPARD using Algorithm 1:
[…]
@@
3.3 $\overline{\mathrm{\textbf{L}}}$EOPARD
The optimization procedure for $\overline{\mathrm{L}}$EOPARD is outlined in Algorithm 1. The goal is to learn a rank-$r$ orthogonal projection $P$ that erases a discrete concept nonlinearly, parameterized via a matrix $U\in\mathbb{R}^{d\times r}$ such that $P=UU^{\top}$, as in Eq. 2. The learning objective to minimize during training combines the projection loss $\mathcal{L}_{p}$ (Eq. 3) and the erasure loss $\mathcal{L}_{\mathrm{MMD}}$ (Eq. 6) and is given by:
@@ 3.3 $\overline{\mathrm{\textbf{L}}}$EOPARD
The number of trainable parameters in $\overline{\mathrm{L}}$EOPARD is $d\cdot r$, significantly lower than that of existing nonlinear erasure methods such as FaRM [3] and KRaM [4], which involve training multi-layer perceptrons for concept removal.
In $\overline{\mathrm{L}}$EOPARD, the adjustable rank $r$ governs the trade-off between the preservation of information and the effectiveness of concept removal. There does not exist a universally optimal value for $r$; the rank adjustability constitutes a key strength of our method, enabling trade-offs between erasure and information preservation that are not possible with more rigid approaches.
Cascaded $\overline{\mathrm{\textbf{L}}}$EOPARD
@@
Cascaded erasure is a two-stage process that integrates linear and nonlinear erasure. In the first stage, we apply an orthogonalized LEACE projection [5] to linearly remove the concept. The embeddings are then projected onto the (dK+1)(d-K+1)-dimensional image subspace of this projection. In the second stage, these linearly "sanitized" embeddings are further processed using a nonlinear erasure method. When instantiated with $\overline{\mathrm{L}}$EOPARD, it amounts to computing two orthogonal projections acting on nested subspaces. The overall cascaded projection remains orthogonal and inherits the benefits of both linear and nonlinear concept erasure. The full procedure for Cascaded $\overline{\mathrm{L}}$EOPARD is outlined in Algorithm 2.
@@ 4.1 Datasets
is a real world dataset suited for studying gender biases in biography classification tasks [10]. It consists in biographies collected through web scraping and labeled with binary gender and occupation (28 in total). We use the version from [23], which contains $\sim$98% of the original dataset, as the full version is no longer publicly available. It consists of 399,423 biographies stratified by occupation, split 65%/10%/25% across training, validation, and test.
@@ 4.2 Practical Aspects of MMD Evaluation
[…]
@@ 4.3 Evaluation metrics
When there is no predefined downstream task for which we can evaluate the accuracy, to measure the overall preservation of the semantic information in the embedding space we rely on the metric $A_{k}$ introduced in [4] that quantifies the average overlap between the $k$-nearest neighbor sets of $x_{i}$ and $x_{i}^{\perp}$. Formally,
$A_{k}=\frac{1}{N}\sum_{x\in\mathcal{D}}\left[\frac{1}{k}|k\mathrm{nn}(x)\cap k\mathrm{nn}(x^{\perp})|\right]\,$, (9)
where $k\mathrm{nn}(\cdot)$ computes the $k$-nearest neighbor set of a representation. The metric ranges from 0 to $1$, with values closer to $1$ indicating greater preservation of the local geometry of the embeddings. We set $k=\frac{|\mathcal{D}|}{2}$, denoting the average neighborhood overlap as $A_{50\%}$.
@@
$\mathrm{DP}=\sum_{y\in\mathcal{Y}}|p(\hat{Y}=y|Z=c)-p(\hat{Y}=y|Z=c^{\prime})|\\$ (11)
@@ 4.4 Implementation Details
All models are trained using the Adam optimizer [17]. For $\overline{\mathrm{L}}$EOPARD, weight decay is set to zero to avoid interfering with the optimization of the projection loss (Eq. 3).
@@ 5.1 Nonlinear Concept Erasure
$\overline{\mathrm{L}}$EOPARD enables effective erasure of information associated with a discrete concept, with the projector rank $r$ serving as a tunable parameter that controls the strength of the erasure. Empirically, we observe that the difficulty of recovering the concept post-erasure increases steadily as the number of preserved dimensions decreases (see Figure 3, left panels). In Table 1, we exhibit configurations where $\overline{\mathrm{L}}$EOPARD outperforms nonlinear erasure baselines in terms of MDL or probing accuracy which drops to near-random levels. Notably, in Table 1(a), gender treated as a ternary concept during training and evaluation is almost fully erased, demonstrating the validity of $\overline{\mathrm{L}}$EOPARD beyond binary settings. Qualitatively, the visualizations in figure 2 and in figures 6,7 in Supplementary Material D illustrate that concept class groups, which were previously distinguishable, become indistinguishable following erasure via $\overline{\mathrm{L}}$EOPARD.
Figure 2: t-SNE visualization of GloVe embeddings before (left) and after (right) erasure via Cascaded $\overline{\mathrm{L}}$EOPARD with a rank-150 orthogonal projection. Gender is a ternary concept. Male-biased, female-biased, and neutral words are displayed in different colors.
Results reported in the left panels of Figure 3 highlights the superior erasure capabilities of Cascaded $\overline{\mathrm{L}}$EOPARD (Algorithm 2). For a fixed projection rank, the difficulty of recovering the concept post-erasure is consistently higher with Cascaded $\overline{\mathrm{L}}$EOPARD compared to the standard training (Algorithm 1). For a given value of MDL, Cascaded $\overline{\mathrm{L}}$EOPARD retains a larger number of dimensions in the feature space. This confirms our intuition that $\overline{\mathrm{L}}$EOPARD trained in a standard way struggles with the removal of strong linear signals, which undermines nonlinear erasure performance. However, concept erasure is achieved with standard $\overline{\mathrm{L}}$EOPARD at low rank. Both FaRM and KRaM, in their standard forms, perform well in terms of both linear and nonlinear erasure, suggesting that implementing them in a cascaded setup should not improve their performance, which we confirm experimentally (see Supplementary Material D).
Table 1: Concept erasure evaluation. Accuracies are reported in $\%$ and MDL in kBits.
@@
$a_{z}^{\mathrm{bin}}$ $\downarrow$ MDL $\uparrow$ $a_{z}^{\mathrm{ter}}$ $\downarrow$ $A_{50\%}\uparrow$ WS-353 $\uparrow$
@@
$a_{z}$ $\downarrow$ MDL $\uparrow$ $a_{y}$ $\uparrow$
@@
$a_{z}$ $\downarrow$ MDL $\uparrow$ $a_{y}$ $\uparrow$
@@
The right panels of Figure 3 highlight a consistent trade-off between concept erasure and the utility of post-erasure representations across all benchmark datasets, with optimal configurations located in the upper-right regions. The rank $r$, which controls the strength of erasure, modulates this trade-off. FaRM and KRaM achieve strong erasure but at the cost of substantial degradation in utility for gender, as evidenced by their positions in the upper-left regions of the left panels in Figures 3(a) and 3(b).
The adjustable rank of $\overline{\mathrm{L}}$EOPARD offers flexibility in navigating this trade-off, enabling configurations that outperform baselines across several settings. Notably, on GloVe embeddings, $\overline{\mathrm{L}}$EOPARD better preserves both local neighborhood structure and semantic content (Table 1(a)).
On Bias in Bios, FaRM and KRaM outperform $\overline{\mathrm{L}}$EOPARD in terms of MDL-based erasure at the expense of downstream task performance. At low rank, $\overline{\mathrm{L}}$EOPARD achieves trade-offs between probing accuracy and downstream task accuracy comparable to baselines (e.g., $r=10$ and $r=50$ for standard and cascaded training, respectively, in Table 1(b)). Figure 3(b) reveals that beyond a certain threshold, performance plateaus; further reducing $r$ yields marginal gains in erasure while significantly harming utility. In such cases, $\overline{\mathrm{L}}$EOPARD (e.g., at $r=150$ with cascaded training) allows for a more satisfying balance between the reduction of the probing accuracy and downstream accuracy compared to baselines.
On DIAL, for race erasure, $\overline{\mathrm{L}}$EOPARD remains competitive without outperforming the baselines. Relying on linear projections in $\overline{\mathrm{L}}$EOPARD provides advantages in simplicity, interpretability, and parameter efficiency. However, such linear methods are inherently limited when addressing the erasure of concepts encoded nonlinearly. While erasing even a small number of dimensions can significantly reduce the recoverability of the target concept, achieving strong erasure often requires neutralizing a large subspace, which may degrade the utility of the resulting representations.